Suppose that we have two events, A and B, with P(A) = 0.50, P(B) = 0.60, and P(A ∩ B) = 0.05. If needed, round your answer to three decimal digits.

(a) Find P(A | B).

(b) Find P(B | A).

(c) Are A and B independent? Why or why not? A and B _____ independent, because _____ P(A).
Answer #1

Solution:

a)

P(A | B) = P(A B ) / P(B)

= 0.05 / 0.60

= 0.083

b)

P(B | A) = P(A B ) / P(A)

= 0.05 / 0.50

= 0.100

c)

P(A B ) = P(A) * P(B)

= 0.50 * 0.60

= 0.300

A and B are not independent, because the probability of A and B is not same.